OpenStudy (anonymous):

how do you find the equation for a graph with x-intercepts of (2,0) and (-6,0) and a maximum point of (-2,4)

OpenStudy (amistre64):

thats a quadratic

OpenStudy (amistre64):

the most basic approach is to solve for 3 equations with variables of a b and c; using your x and y values as your quadratic constants

OpenStudy (amistre64):

or...or factor in the roots

OpenStudy (amistre64):

(x-2)(x+6)

OpenStudy (amistre64):

C(x^2 +4x -12) is the family of quadratics with these roots

OpenStudy (amistre64):

if we complete the square and plug in for the vertex we got it solved

OpenStudy (amistre64):

x^2 +4x +4 = 12+4 (x+2)^2 = 16

OpenStudy (amistre64):

(-1/4)(x+2)^2 +4 = y might be it

OpenStudy (amistre64):

lets check it out :)

OpenStudy (amistre64):

-1/4(x^2 +4x-12) -x^2/4 -x +3 1/2(-1/4) = 1/(-2/4) = 1/(-1/2) = -2 x is good

OpenStudy (amistre64):

-1/4(x^2 +4x -12) |x=-2 (-1/4)(4 -8 -12) = -16/-4 = 4 YES

OpenStudy (amistre64):

now that I taught mysef something, any questions lol

OpenStudy (anonymous):

no

OpenStudy (amistre64):

:) good, cause I dont know if I can remember how I did it ;)